1. Consider the reaction:
5 A + 3 B →→ 2 C

The rate of the reaction is found to be 0.0486 M/s. What is the rate of change of B? Be sure to consider whether the substance is disappearing or appearing in your answer.

2. For the following zero order reaction, the rate constant is 0.0311. If [A]0 = 8.68, calculate [A] at 2.9 seconds. 2 A →→ 3 B
Answer #1

1. Rate = -1/3 de dB = - 3x0.0486 M 07 =-0.1458 M he sign represents disappearance. 20 R = Ro -kt Ca] = [Alo - kt ( 24 → 38)